• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name lkaufman

    Log entry time 17:41:02 on September 01, 2005

    Entry number 154157

    keyword=Low I cavity commis. part 3

    We lowered the current to ~ 50 nA, and Arne Freyberger set up the
    cavities for data taking. Raster is off for these runs.

    Spectr. DAQ HAPPEX DAQ Comment
    2603 4918 box scan
    2604 4919 box scan, different post-gains

    Back to production/ready for linearity study.